Maria, an experienced shipping clerk, can fill a certain order in 6 hours. Bill, a new clerk, needs 11 hours to do the same job. Working together, how long will it take them to fill the order?
Answer by scott8148(6628)

their combined time determines the fraction of the job each one does
___ the fractions add to one (the whole job)

(x / 6) + (x / 11) = 1

11x + 6x = 66

x = 66 / 16 = 4.125